OpenStudy (anonymous):

William folds his origami paper along the dotted lines shown below so that the triangular flaps overlap exactly on the square. If the dimensions of the origami paper are decreased by One-half, what will be the new area of the paper?

OpenStudy (anonymous):

5 square inches 20 square inches 25 square inches 10 square inches

OpenStudy (cwrw238):

the area of origami paper before decrease = 100 square inches if dimnessions decrease by 1/2 then area decreases by 1/2 ^ 2 = 1/4 so new area = 100 8 1/4 = 25 sq ins

OpenStudy (cwrw238):

* 100 * 1/4

OpenStudy (anonymous):

got it thanks @cwrw238

OpenStudy (cwrw238):

your welcome